[Brahms-l] Shift report 20040217 08:00-16:00

From: Apache <apache@bnl.gov>
Date: Tue Feb 17 2004 - 15:37:50 EST
------------------------------------------------------------
Date: Tue Feb 17 15:37:50 2004, Shift: 08:00-16:00
Supervisor: Ionut ARSENE, with: 
Log book#15 pages 60 - 60

Shift Summary:
------------------------------------------------------------
1. We had only one hour of beam, but we couldn't take data 
because of RHIC clock problems. Our ZDC rates are ~150, and 
trig2,3 and 6 rates 0. Dana tried to solve the problem. Our 
magnets tripped while Dana tried to reboot the VME crate 
(i think) and caused (again) a RHIC Quench Link Interlock.
The magnets are OK now.

2. No achievemets due to the above problems, except pedestals.
 10408 - tpc pedestals
 10409 - non-tpc pedestals

3. No detector problems observed.  
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue Feb 17 15:38:03 2004

This archive was generated by hypermail 2.1.8 : Tue Feb 17 2004 - 15:38:24 EST